# Neural network embeddings recover value dimensions from psychometric survey items on par with human data

### v1

@@ 4 Results
The results presented here are derived from SQuID treated embeddings from the best-performing model "Linq-Embed-Mistral".
@@ 4.1.1 Internal consistency
Following Schwartz and Cieciuch (2022), we employ Cronbach’s alpha to assess the internal consistency of latent dimensions. Through functional equivalence in this context (for details see Subsection A.3), we can directly compare embeddings of all our candidate models with human data [^3] We retrieve the human values from Table S2 in SI of Schwartz and Cieciuch (2022) which reports averages over 49 countries.. Cronbach’s alpha compares the amount of shared variance among the items of a dimension to the amount of overall variance. In previous research, performance baselines of random embeddings have been shown to be higher than intuitively expected in some scenarios Wieting and Kiela (2019). To assess which baseline levels we reach, we produce a vector with 4096 normally distributed values ($M_{ij}\sim\mathcal{N}(0,\,1)$) for each item. We repeat the procedure 10 000 times and report average values of Cronbach’s alpha for each dimension in these random embeddings. In addition to reporting all 19 values for each model, human data and random embeddings, we also provide an overall average over all dimensions. In Table 1, we find that Linq-Embed-Mistral is the best performing approach. It beats the other approaches 14 out of 19 times and has the highest average Cronbach’s alpha over all dimensions (including being slightly higher than the human data).
@@ 4.1.2 Dimension-dimension correlations of human rater judgments and embeddings
We create a dimension-dimension similarity matrix using Pearson correlation for both the embeddings (shown in Subsection A.6) and the human data. The 171 pairs of both types of data positively correlate substantially: $0.74$ [$0.66,0.80$]. Figure 1 shows the scatter plot and parameters of a simple linear model that explains $55\%$ of the variance observed.
Figure 1: Relationship between item similarity (correlation) derived from embeddings (SQuID treated) and human data. We explain more than half of the variance with a simple linear model.
@@ 4.2 MDS Configuration from Embeddings
We turn the dimension-dimension correlation matrix of Subsection A.6 into a dissimilarity matrix. We use multidimensional scaling444mds with ’type=”ordinal”’ from the R package smacof Mair et al. (2022), for details see code file. to reduce the high-dimensional dissimilarity to a two dimensional space that preserves distances between dimensions. Figure 5 shows the resulting circular structure.
4.3 Procrustes Matrix Transformation
@@
Next, we want to assess quantitatively the correspondence of the embeddings MDS configuration to the empirical configuration of Schwartz and Cieciuch (2022) [^5] We compute the human MDS configuration from the pooled correlation matrix over all 49 countries shared by the authors as Table S10 of SI. derived from human rater judgments. As MDS configurations are not unique, arbitrary matrix transformations such as translation, rotation and reflection are admissible because these operations do not change pairwise distances. We use Procrustes matrix transformation666procrustes from the R package vegan with the human configuration as target and the embeddings configuration as testee. This way, we transform the embeddings configuration to achieve maximum similarity with the human rater judgments configuration.
4.4 Comparison of Configurations from Human Data and from Embeddings
To quantitatively assess the similarity of the two MDS configurations after Procrustes transformation of the embeddings MDS configuration matrix, we compute cosine similarities of the vectors of the two first dimensions (X coordinates in the configuration plot) as well as the two second dimensions (Y coordinates). These congruences coefficients are $0.88$ for the first and $0.82$ for the second dimension. To further assess similarities and differences between individual dimensions, Figure 2 allows for visual assessment of the two configurations.
@@ 4.4 Comparison of Configurations from Human Data and from Embeddings
5 Discussion
Our validation checks uncover that value dimensions contained in neural embeddings can display internal consistency on par with those derived from human data (or even above). As a comparison baseline, we find no meaningful values for randomly initiated embeddings. Correlation of dimension-dimension similarities for human data and embeddings show substantial agreement between the two types of data. We find no bias in the regressing one on the other, a strong positive relationship and a moderately high coefficient of determination for a simple linear model.
In the embeddings configuration of Figure 5, we can already see the circular structure emerging that is postulated by Schwartz and Cieciuch (2022). Values that are facets of the same phenomenon (like "Universalism-Concern", "Universalism-Nature" and "Universalism-Tolerance") are placed close together. Values that are seen as distant to Universalism (such as "Power Dominance" and "Power Resource") are placed far apart on the circular structure, roughly on the opposite side. The same can be observed with the facets of "Self-Direction" in comparison to the facets of "Conformity". The results allow for a visual interpretation that follows quite closely along the lines of the canonical theory of value orientations that powers the PVQ-RR.
Comparing the two MDS configuration after Procrustes transformation shows fair similarity according to conventional interpretations of congruence coefficients (for details see Section A.5). That we find the two configurations to be fairly similar on this formal criterion is quite surprising given that they derive from two very distinct data sources. Inspection of individual value dimensions in Figure 2 reveals strong overlaps of the two solution with the facets of Security (Security Personal and Security Societal) and basically identical placement of some dimensions such as Conformity-Interpersonal, Power Resources and Universalism-Tolerance. While other are placed closely like Self-direction Thought and Power Dominance, dimensions can also be distinctly different between the two configurations, for example Stimulation (ST) that is placed much closer to Hedonism (HE) in the embeddings configuration.
@@ 6 Conclusion
We introduced our methodological advancement of SQuID and show with the example of the PVQ-RR that it allows an embedding approach to reach performance on par with human data in extracting latent value dimensions according to several evaluation metrics.
@@
We envision many new directions of research to emerge from psychometric approaches to AI. As one example, our approach allows for a systematic and rigorous investigation of potential cultural biases contained in LLMs: Schwartz and Cieciuch (2022) confirmed their theories of value orientations with data from 49 different cultures using translated questionnaires. We see an exciting opportunity to use our approach and MDS configurations on the level of those individual countries to systematically assess closeness or distance of an LLM to certain cultural value orientations quantitatively by computing congruence coefficients and other error metrics. Generally, working on the embedding level as we propose is limiting sources of potential methodological biases stemming from token extraction strategies that are necessary when working with generated text instead.
@@
We view our proposal as complimentary to both traditional workflows as well as to novel approaches in the (social) sciences. By adopting embeddings approaches, we don’t aim to replace human participants. Also well into the future, researchers will have very good reasons to work with humans, at the bare minimum to provide the necessary reference points for comparison as we demonstrated in our analyses. But different to ideas discussed before, such as "pseudo-factor analysis" Guenole et al. (2025), we want to go beyond merely creating another step in the typical validation pipeline in psychometrics. We see potential beyond that: With these novel technologies we can potentially investigate the full and diverse spectrum of expression in all natural languages. This creates novel opportunities to expand the scope of human behavior and experience represented in measurement tools.
@@
As mentioned in several places in the manuscript, future research will have to uncover more of the peculiarities of that novel kind of data in the form of semantic embeddings for (social) scientific applications. More experience with the data will in turn help the methods (and the theories). We see "neural embeddings as human rater judgments" as a good working hypothesis and don’t want to imply that both data are exactly the same and capture human behavior and experience in exactly the same way. As with other synthetic data, novel treatments such as specific transformation (in addition to general methodologies like SQuID) and other processing steps may be necessary depending on the application context.
@@
The notion of "similarity" is actually not well-defined as a classical discussion in philosophy shows. The argument goes against the existence of a strong logical underpinning of notions of similarity, while it’s practical relevance is acknowledged. Similarity is considered "still serviceable in the streets" Goodman (1972). Some parts of the original logical argument have been debated and countered Leitgeb (2007). We further want to note that our metaphor of semantic embeddings as human rater judgments helps us to prevent some concerns: Each dimension of semantic embeddings as the equivalent to a human rater judgment provides the necessary context ("similar to what?") to justify relative assessments of similarity for our use case.
@@
Still, crowdworking delivers incomes to people around the world that come from many marginalized groups. Although it can leave workers underpaid, it is an integral and existential part of many lives. Failures of platforms to react and dwindling numbers of crowdworking jobs (that are instead outsourced to LLMs) can leave many people worse off than before.
On another, more distant future, level, we could also see some danger in people accepting analyses based on semantical embeddings as some kind of "gold standard", a LLM normality to compare with the results of people, forgetting about many possible biases. Abstractly, like any assessment, this could increase the danger of diagnosing somebody as abnormal vis-a-vis an unattainable standard.
Generally, we see potential harm to fall mostly proportional on all parts of the population in terms of race, gender or class. However, the lack of jobs for crowdworkers can be expected to have a disproportional affect on people on low incomes in less affluent (and predominantly English-speaking) parts of the world. In addition, we try to limit the potential harm of creating novel synthetic LLM norms by restricting ourselves to a questionnaire of value orientations such as the PVQ-RR that is rather uncontroversial and is not positing any "correct" or "normal" value orientations for individuals.
@@ Appendix A Appendix
In our paper repository [^7] https://anonymous.4open.science/r/embeddings_values-BFBE/, we share documented code that allows other researchers to fully replicate our results using openly available data.
@@ A.3 Cronbach’s Alpha of human rater judgments and embeddings
$\alpha={k\over k-1}\left(1-{\sum_{i=1}^{k}\sigma^{2}_{y_{i}}\over\sigma^{2}_{y}}\right)$
@@ A.4 PCA instead of MDS
As a robustness check, replacing MDS with PCA888prcomp in R with ’center = TRUE’ and ’scale = TRUE’ and plotting the first two dimensions shows similar clustering to the MDS solution but less circular structure as PCA does not guarantee to preserve distances:
@@ A.5 Congruence coefficients
Labels refer to Embeddings and human rater judgments from Schwartz Schwartz and Cieciuch (2022). Congruence coefficients are computed for each of the two dimensions after transformation of the embeddings configuration. There is no strict standard, but according to typical interpretations, values in the range of $0.85\text{-}0.94$ between two compared configurations display fair similarity Lorenzo-Seva and Ten Berge (2006). To establish significance, we create 2 times 1000 random vectors, use Procrustes transformation to maximize similarity between them and compute their cosine similarity. We repeat this procedure 1000 times and find $p<0.001$ for the cosine similarities we report above over all runs, indicating that they are significantly different from a random baseline and cannot be explained as an artifact of Procrustes transformation.
@@ A.6 Effect of Questionnaire Embedding Subtraction
Figure 4, from left to right, shows that through SQuID we achieve negative correlations in the item-item similarity matrix (sorted alphabetically by dimensions). The similarity matrix of the same SQuID treated embeddings averaged by dimension:
We compute correlation coefficients between SQuID treated embeddings that are averaged by dimension. We don’t show the diagonal of 1s to better allow to observe correlation ranges.

### v2

@@ 5 Results
Unless otherwise noted, the results presented here are derived from SQuID treated embeddings from the best-performing model "Linq-Embed-Mistral". For performance of all models compared to the human rater judgment data as measured with factor congruence scores, see Table 3.
@@ 5.1.1 Internal consistency
Following Schwartz and Cieciuch (2022), we employ Cronbach’s alpha to assess the internal consistency of latent dimensions. Through functional equivalence in this context (for details see Subsection A.3), we can directly compare embeddings of all our candidate models with human data [^3] We retrieve the human rater judgment data from Table S2 in SI of Schwartz and Cieciuch (2022) which reports averages over 49 countries.. Cronbach’s alpha compares the amount of shared variance among the items of a dimension to the amount of overall variance. In previous research, performance baselines of random embeddings have been shown to be higher than intuitively expected in some scenarios Wieting and Kiela (2019). To assess which baseline levels we reach, we produce a vector with 4096 normally distributed values ($M_{ij}\sim\mathcal{N}(0,\,1)$) for each item. We repeat the procedure 10 000 times and report average values of Cronbach’s alpha for each dimension in these random embeddings. In addition to reporting all 19 values for each model, human data and random embeddings, we also provide an overall average over all dimensions. In Table 1, we find that Linq-Embed-Mistral is the best performing model. It beats the other approaches 14 out of 19 times and has the highest average Cronbach’s alpha over all dimensions (including being slightly higher than the one measured from the human data).
@@ 5.1.2 Dimension-dimension correlations of human rater judgments and embeddings
We create a dimension-dimension similarity matrix (using Pearson correlation) for both the embeddings (shown in Subsection A.6) and the human data. The 171 pairs of both types of data positively correlate substantially: $0.74$ [$0.66,0.80$]. Figure 1 shows the scatter plot and parameters of a simple linear model that explains $55\%$ of the variance observed.
Figure 1: Relationship between item similarity (Pearson correlation) derived from embeddings (SQuID treated) and human data. We explain more than half of the variance with a simple linear model.
@@ 5.2 MDS configuration from embeddings
We turn the dimension-dimension correlation matrix of Subsection A.6 into a dissimilarity matrix. We use multidimensional scaling444mds with ’type=”ordinal”’ from the R package smacof Mair et al. (2022), for details see code file. to reduce the high-dimensional dissimilarity matrix to a two dimensional space that preserves distances between dimensions. Figure 2 shows the resulting circular structure.
5.3 Procrustes matrix transformation to compare human data and embeddings
@@
Next, we want to compare the embeddings MDS configuration to the empirical configuration of Schwartz and Cieciuch (2022) [^5] We compute the human MDS configuration from the pooled correlation matrix over all 49 countries shared by the authors as Table S10 of SI. derived from human rater judgments. As MDS configurations are not unique, arbitrary matrix transformations such as translation, rotation and reflection are admissible because these operations do not change pairwise distances. We use Procrustes matrix transformation666procrustes from the R package vegan with the human configuration as target and the embeddings configuration as testee. This way, we transform the embeddings configuration to achieve maximum similarity with the human rater judgments configuration. To quantitatively assess the similarity of the two MDS configurations after Procrustes transformation of the embeddings MDS configuration matrix, we compute cosine similarities of the vectors of the two first dimensions (X coordinates in the configuration plot) as well as the two second dimensions (Y coordinates). These congruence coefficients are $0.88$ for the first and $0.82$ for the second dimension. To further assess similarities and differences between individual dimensions, Figure 3 allows for visual assessment of the two configurations.
5.4 Generalizability to other questionnaires
To assess whether SQuID generalizes beyond value theory, we apply the method to three widely-used personality questionnaires: the International Personality Item Pool (IPIP) Goldberg (1992), the Big Five Inventory-2 (BFI-2) Soto and John (2017), and the HEXACO Personality Inventory Lee and Ashton (2018). These personality measures differ from the PVQ-RR in both theoretical structure and item content, providing a rigorous test of SQuID’s domain generality. Similar to value theory, a challenge in personality measurement is capturing negative correlations between items with opposing content. The range of inter-item correlations thus serves as a critical metric: wider ranges indicate that embeddings successfully capture both positive associations between similar items and negative relationships between opposing items. Table 2 shows that SQuID substantially increases correlation ranges across all models for the IPIP questionnaire. Raw embeddings exhibit limited ranges (0.14 to 0.82 depending on model), but SQuID treatment expands these dramatically. The average relative gain in range is 261.5%, with individual models showing improvements from 21.20% (mpnet, which was already optimized for personality content through domain-specific finetuning) to 724.39% (kalm). Notably, even models that initially showed reasonable raw ranges (e.g., mpnet at 0.82) benefit from SQuID. We observe similar patterns for BFI-2 (average relative gain over all models: 207.2%) and HEXACO (147.9%). These consistent improvements across multiple personality inventories suggest that SQuID can be applied broadly to psychometric measurement.
Raw Range SQuID Range Abs. Gain Rel. Gain (%)
gemini 0.35 0.92 0.57 161.88
jina 0.66 1.30 0.64 95.97
kalm 0.14 1.19 1.04 724.39
linq 0.23 0.92 0.69 304.13
mpnet 0.82 1.00 0.17 21.20
Table 2: Comparison of raw vs SQuID embeddings across all models for the IPIP personality questionnaire. Raw Range and SQuID Range show correlation ranges. Absolute Gain shows the increase in range, while Relative Gain shows the percentage improvement.
6 Discussion
Our validation checks uncover that value dimensions contained in neural embeddings can display internal consistency on par with human data (or even above). As a comparison baseline, we find no meaningful values for randomly initiated embeddings. Correlation of dimension-dimension similarities for human data and embeddings show substantial agreement between the two types of data. We find no bias in regressing one on the other, a strong positive relationship and a moderately high coefficient of determination for a simple linear model.
In the embeddings configuration of Figure 2, we can see the circular structure emerging that is postulated by Schwartz and Cieciuch (2022). Values that are facets of the same phenomenon (like "Universalism-Concern", "Universalism-Nature" and "Universalism-Tolerance") are placed close together. Values that are seen as distant to Universalism (such as "Power Dominance" and "Power Resource") are placed far apart on the circular structure, roughly on the opposite side. The same can be observed with the facets of "Self-Direction" in comparison to the facets of "Conformity". The results allow for a visual interpretation that follows quite closely along the lines of the canonical theory of value orientations that powers the PVQ-RR.
Comparing the two MDS configurations after Procrustes transformation shows fair similarity according to conventional interpretations of congruence coefficients (for details see Section A.5). That we find the two configurations to be fairly similar on this formal criterion is quite surprising given that they derive from two very distinct data sources. Inspection of individual dimensions in Figure 3 reveals strong overlaps (e.g., Security facets, Conformity-Interpersonal, Power Resources) and some differences (e.g., Stimulation is placed closer to Hedonism in the embeddings configuration).
@@ 7 Conclusion
We introduced our methodological advancement of SQuID and show with the example of the PVQ-RR that it allows an embedding approach to reach performance on par with human data in extracting latent value dimensions according to several evaluation metrics. We envision many new directions of research to emerge from psychometric approaches to AI. As one example, our approach allows for a systematic and rigorous investigation of potential cultural biases contained in LLMs: Schwartz and Cieciuch (2022) confirmed their theories across 49 countries using translated questionnaires. We see an exciting opportunity to use our approach and MDS configurations on the level of those individual countries to systematically assess closeness or distance of an LLM to certain cultural value orientations quantitatively by computing congruence coefficients and other error metrics. Working on the embedding level, as we propose, limits sources of potential methodological biases like token extraction strategies that are necessary when working with generated text instead.
The fundamental difference between human rater judgments and neural embeddings as data sources raises important questions about what each captures. Human ratings reflect conscious, individual self-assessment assumed to be influenced by the traits to be measured. In contrast, embeddings represent semantic relationships derived from statistical regularities in language use learned from vast text corpora by neural models. That these two data sources converge to produce similar psychometric structures (R² of 0.55, fair congruence coefficients) suggests that they nevertheless tap into shared underlying semantic spaces. The 45% of variance not explained by our linear model could stem from genuine differences in what the two data types measure, psychometric noise in human responses (e.g. response styles, measurement error) or systematic biases in language model training data. Future research should decompose this unexplained variance to better understand the complementary information each data source provides.
While our results demonstrate SQuID’s effectiveness for value orientations (PVQ-RR) and initial tests show promise for personality inventories (IPIP, BFI-2, HEXACO), systematic evaluation across additional psychometric domains remains an important direction for future research. We suggest several priority areas: First, researchers should test SQuID on questionnaires with different structural properties. The PVQ-RR exhibits a circumplex structure suited to multidimensional scaling, while personality inventories typically employ factor-analytic approaches. Testing on hierarchical models, formative constructs or network-based psychometric structures would clarify boundary conditions. Second, questionnaires employing reverse-keyed items require methodological refinement. As noted in our limitations, directly "reversing" embeddings is conceptually unclear. One possible approach could involve computing separate scale embeddings for regular and reverse-keyed items, then subtracting separately from the two types of items. This extension would at least better handle the intentional polarity inversions common in personality and attitude measures. Third, extend testing to attitudes, motivational orientations and clinical inventories, which may require methodological adaptations. Finally, multilingual testing would establish whether SQuID’s benefits generalize across languages.
@@
We view our methodological proposal as complementary to both traditional workflows as well as to novel approaches in the (social) sciences. By adopting embeddings approaches, we don’t aim to replace human participants. Also well into the future, researchers will have good reasons to study humans, at the bare minimum to provide necessary reference points for comparison as we demonstrated in our analyses. But different to ideas discussed before, such as "pseudo-factor analysis" Guenole et al. (2025), we want to go beyond merely creating another step in the typical validation pipeline in psychometrics. We see potential beyond that: With these novel technologies we can potentially investigate the full and diverse spectrum of expression in all natural languages. This creates novel opportunities to expand the scope of human behavior and experience represented in measurement tools.
@@
Future research will have to uncover more of the peculiarities of that novel kind of data in the form of semantic embeddings for (social) scientific applications. More experience with the data will in turn help the methods (and the theories). We see "neural embeddings as human rater judgments" as a good working hypothesis and don’t want to imply that both data are exactly the same and capture human behavior and experience in exactly the same way. As with other synthetic data, novel treatments such as specific transformation (in addition to general methodologies like SQuID) and other processing steps may be necessary depending on the application context.
@@
The notion of "similarity" is actually not well-defined as a classical discussion in philosophy shows. The argument goes against the existence of a strong logical underpinning of notions of similarity, while its practical relevance is acknowledged. Similarity is considered "still serviceable in the streets" Goodman (1972). Some parts of the original logical argument have been debated and countered Leitgeb (2007). We further want to note that our metaphor of semantic embeddings as human rater judgments helps us to prevent some concerns: Each dimension of semantic embeddings as the equivalent to a human rater judgment provides the necessary context ("similar to what?") to justify relative assessments of similarity for our use case.
@@
Still, crowdworking delivers income to people around the world that come from many marginalized groups. Although it can leave workers underpaid, it is an integral and existential part of many lives. Failures of platforms to react and dwindling numbers of crowdworking jobs (that are instead outsourced to LLMs) can leave many people worse off than before.
On another, more distant future, level, we could also see some danger in people accepting analyses based on semantical embeddings as some kind of "gold standard", a LLM normality to compare with the results of humans, forgetting about many possible biases. Abstractly, like any assessment, this could increase the danger of diagnosing somebody as abnormal vis-a-vis an unattainable standard.
Generally, we see potential harm to fall mostly proportional on all parts of the population in terms of race, gender or class. However, potential lack of jobs for crowdworkers can be expected to have a disproportionate effect on people on low incomes in less affluent (and predominantly English-speaking) parts of the world. In addition, we try to limit the potential harm of creating novel synthetic LLM norms by restricting ourselves to a questionnaire of value orientations such as the PVQ-RR that is rather uncontroversial and is not positing any "correct" or "normal" value orientations for individuals.
@@ Appendix A Appendix
In our paper repository [^7] https://github.com/maxpel/embeddings_values, we share documented code that allows other researchers to fully replicate our results using openly available data.
@@ A.3 Cronbach’s Alpha of human rater judgments and embeddings
$\alpha={\frac{k}{k-1}}\left(1-{\frac{\sum_{i=1}^{k}\sigma^{2}_{y_{i}}}{\sigma^{2}_{y}}}\right)$
@@ A.4 PCA instead of MDS
As a robustness check, we replace MDS with PCA888prcomp in R with ’center = TRUE’ and ’scale = TRUE’. In Figure 5, we plot the first two dimensions that show similar clustering to the MDS solution but less circular structure since PCA does not guarantee to preserve distances.
Figure 5: Replacing MDS with PCA. We plot the first two dimensions found through PCA with the embeddings. We see similar clustering as compared to the MDS configuration with less overall circular structure.
Figure 6: Congruence coefficients used to measure the similarity between our two MDS configurations. Labels refer to Embeddings from linq and human rater judgments from Schwartz Schwartz and Cieciuch (2022).
@@ A.5 Congruence coefficients
Congruence coefficients are computed for each of the two dimensions after Procrustes transformation of the embeddings configuration. There is no strict standard, but according to typical interpretations, values in the range of $0.85\text{-}0.94$ between two configurations indicate fair similarity Lorenzo-Seva and Ten Berge (2006). To test significance, we create pairs of 1000 random vectors, use Procrustes transformation to maximize their similarity and compute cosine similarity. We repeat this procedure 1000 times and find $p<0.001$ for the cosine similarities we report over all runs, indicating that they are significantly different from random baselines and cannot be explained as artifacts of Procrustes transformation.
Table 3 presents these scores for all five models tested. For matching dimensions (SD1_ED1 and SD2_ED2), higher congruence coefficients indicate better alignment with the human-derived structure. For non-matching dimensions (SD1_ED2 and SD2_ED1), lower values indicate proper orthogonality. Consistent with our other evaluation measures in the main text such as Cronbach’s Alpha, linq achieves the strongest performance with congruence scores of 0.88 and 0.82 for the two matching dimensions, substantially outperforming other approaches. Gemini also performs quite well, while the other models show weaker congruence.
SD1 ED1 SD1 ED2 SD2 ED1 SD2 ED2
gemini 0.72 0.05 0.05 0.78
jina 0.37 0.24 0.26 0.78
kalm 0.61 0.13 0.14 0.57
linq 0.88 0.05 0.05 0.82
mpnet 0.53 0.29 0.31 0.26
Table 3: Factor congruence scores with human rater judgments for all models. For matching dimensions (SD1_ED1 and SD2_ED2) higher is better. For differing dimensions (SD1_ED2 and SD2_ED1) lower is better.
@@ A.6 Effect of Questionnaire Embedding Subtraction
Figure 8: Similarity matrix of SQuID treated embeddings averaged by dimension. We compute correlation coefficients between SQuID treated embeddings that are averaged by dimension. We don’t show the diagonal of 1s to better allow to observe correlation ranges.
Figure 7, from left to right, shows visually that with SQuID we achieve negative correlations in the item-item matrix (dimensions sorted alphabetically).